In each of the following cases, calculate the accounting break-even and the cash break-even points. Ignore any tax effects in calculating the cash break-even. (Do not round intermediate calculations and round your answers to 2 decimal places, e.g., 32.16.)

Case Unit Price

Unit

Variable

Cost

Fixed Costs Depreciation
1 $3,040 $2,455 $9,060,000 $3,400,000
2 91 59 67,000 240,000
3 20 5 2,600 960
Case Accounting break-even Cash break-even
1
2
3
Answer #1

Accounting break-even=(Fixed costs+Depreciation)/(Unit price-Unit Variable Cost)

Cash break-even=Fixed costs/(Unit price-Unit Variable Cost)

Case Accounting break-even Cash break-even
1 (9,060,000+3,400,000)/(3,040-2,455)=21299.15(Approx). 9,060,000/(3,040-2,455)=15487.18(Approx).
2 (67,000+240,000)/(91-59)=9593.75 67,000/(91-59)=2093.75
3 (2600+960)/(20-5)=237.33(Approx). 2600/(20-5)=173.33(Approx).